# CitDet: A Benchmark Dataset for Citrus Fruit Detection

### v2

@@ VI-B Discussion
Contrary to other fruit detection works, we found that processing the image in a tiled manner performs better than detectors that operate on the full image. We hypothesize that there are two primary factors for this. The first factor is the difference between the tiling methods used. For example, in the Bargoti and Underwoods [16] model, training images are split into overlapping chunks, which requires the filtering of overlapping bounding boxes. Our tiling method does not have any overlap. Instead, we process the bounding boxes on the edges of the fruits before training and therefore do not require the filtering step at the end. The second factor contributing to the improved performance of the tiled versions is the ability to preserve information when resizing the image for the model.
In comparing the state-of-the-art object detection algorithms, we found that YOLOv7 outperforms all the other algorithms on both the whole and tiled images. However, Faster R-CNN performs the best for small objects on both image types. The results in Tables II and III suggest that the YOLO architectures benefit more from transfer learning on high-resolution images when compared to Faster R-CNN. All the networks demonstrate the capability to distinguish between fruit yield and drop even in ambiguous cases such as in Fig. 3 and 4. We observed that all the methods struggled to accurately detect small object instances. We note that standard convolutional approaches performed better than transformers for small object detection. Nonetheless, future research is needed to accurately detect small and medium sized objects for all methods. Our findings confirm previous works (e.g., [52, 17]), which found that object size is one of the main determinants of error in object detection.
@@ VI-C Yield Estimation
For all experiments we consider only the fruit on the trees for prediction of the fruit yield. We evaluated the per-tree yield prediction correlation with the ground-truth estimates using the square of the Pearson correlation coefficient,
@@
[…]
@@
An analysis of large prediction errors showed that many images contained parts of neighboring trees in the image, which attributed to higher fruit counts. The second method, filter-detect-count, aims to address this issue by using a two-stage detection approach to filter out neighboring trees. The first stage uses a separate model to detect all trees and their canopies. Then, the central most detection was taken as input to the second stage and passed to a YOLO model for fruit detection. The improved method correctly predicted four out of the five top-producing trees with an error of only -1.49% on the total yield estimation of the orchard as shown in Fig. 5.
@@ VII Conclusion and Future Work
In this work we introduced CitDet, a new dataset for citrus detection in orchards prone to citrus greening. The dataset was created with the intention of being diverse and unbiased, it contains many different varieties of citrus at different stages of maturity and health. By distinguishing between fruit on the ground and fruit on the trees, we showed that it is possible to identify fruit that contributes to yield via a single model. With this annotated collection of citrus, we aim to increase the comparability of fruit detection algorithms and advance research on this topic.
Detectors trained on CitDet can be applied for citrus tracking and yield estimation, size estimation, and automated fruit picking in follow-up studies. The results of the object detection benchmarks indicate that CitDet is challenging for modern detectors and there are many opportunities for subsequent research and improvement. In particular, future work is necessary to improve the accuracy of detectors on small objects in cluttered environments.
@@ Acknowledgments
This material is based upon work supported by the United States Department of Agriculture (USDA) under USDA-ARS CRIS #6034-21000-018-000-D, #6066-21310-006-000-D, and USDA-ARS Non-Assistance Cooperative Agreement #6066-21310-005-061-S. The authors thank members of the Hulse-Kemp Lab including Dr. Keo Corak, Dr. Emily Delorean, Ashley Schoonmaker, Grant Billings, and Cassie Newman for assistance in image annotation and USDA-ARS USHRL technicians Jefferson Shaw and Darren Cole for assistance in image collection.

### v3

@@ VI-B Discussion
Contrary to other fruit detection works, we found that processing the image in a tiled manner performs better than detectors that operate on the full image. We hypothesize that there are two primary factors for this. The first factor is the difference between the tiling methods used. For example, in the Bargoti and Underwood [16] model, training images are split into overlapping chunks, which requires the filtering of overlapping bounding boxes. Our tiling method does not have any overlap. Instead, we process the bounding boxes on the edges of the fruits before training and therefore do not require the filtering step at the end. The second factor contributing to the improved performance of the tiled versions is the ability to preserve information when resizing the image for the model.
In comparing the state-of-the-art object detection algorithms, we found that YOLOv7 outperformed all the other algorithms on both the whole and tiled images. However, Faster R-CNN performed the best for small objects on both image types. The results in Tables II and III suggest that the YOLO architectures benefit more from transfer learning on high-resolution images when compared to Faster R-CNN. All the networks demonstrate the capability to distinguish between fruit yield and drop even in ambiguous cases such as in Fig. 5 and Fig. 6. We observed that all the methods struggled to accurately detect small object instances. We note that standard convolutional approaches performed better than transformers for small object detection. Nonetheless, future research is needed to accurately detect small and medium sized objects for all methods. Our findings confirm previous works (e.g., [53, 18]), which found that object size is one of the main determinants of error in object detection.
@@ VI-C Yield Estimation
For all the experiments we consider only the fruit on the trees for prediction of the fruit yield. We evaluated the per-tree yield prediction correlation with the ground-truth estimates using the square of the Pearson correlation coefficient,
@@
where $n$ is the total number of trees in the experiment, $x$ and $y$ are the predicted and ground-truth yield estimates, and $\bar{x}$ and $\bar{y}$ are the mean of the predicted and ground-truth yield estimates. Two simple yield estimation methods were compared: detect-count and filter-detect-count. The R2 values are reported in Table IV.
@@
An analysis of large prediction errors showed that many images contained parts of neighboring trees in the image, which attributed to higher fruit counts. The second method, filter-detect-count, aims to address this issue by using a two-stage detection approach to filter out neighboring trees. The first stage uses a separate model to detect all trees and their canopies. Then, the central most detection is taken as input to the second stage and passed to a YOLO model for fruit detection. The improved method correctly predicted four out of the five top-producing trees with an error of only -1.49% on the total yield estimation of the orchard as shown in Fig. 7.
@@ VII Conclusion and Future Work
In this work we introduced CitDet, a new dataset for citrus detection in orchards prone to citrus greening. By distinguishing between fruit on the ground and fruit on the trees, we showed that it is possible to identify fruit that contributes to yield via a single model. With this annotated collection of citrus, we aim to increase the comparability of fruit detection algorithms and advance research on this topic. Importantly, the metadata and associated tree-level yield estimates are included to serve as a ground-truth metric.
Detectors trained on CitDet can be applied for citrus tracking and yield estimation, size estimation, and automated fruit picking in follow-up studies. The results of the object detection benchmarks indicate that CitDet is challenging for modern detectors and there are many opportunities for subsequent research and improvement. In particular, future work is necessary to improve the accuracy of detectors on small objects in cluttered environments and the inclusion of biological ground-truth data will allow for further analysis.
@@ Acknowledgments
The authors thank members of the Hulse-Kemp Lab including Dr. Keo Corak, Dr. Emily Delorean, Ashley Schoonmaker, Grant Billings, and Cassie Newman for assistance in image annotation and USDA-ARS USHRL technicians Jefferson Shaw and Darren Cole for assistance in image collection.